OpenStudy (firejay5):

Using KE = 1/2mv^2, solve: What is the mass of a baseball that has a kinetic energy of 105J and is traveling at 10m/s? *Is it still possible to use that equation to get the answer of 2.1kg and if so please solve it using that equation.*

OpenStudy (firejay5):

@confluxepic

OpenStudy (anonymous):

Yes, it is infact possible. Using the formula you mentioned, you just have to transpose and make m the subject of the formula. Do you know how to transpose?

OpenStudy (anonymous):

In other words, you just have to rearrange the formula.

OpenStudy (firejay5):

I need it worked out correctly and accurately @kym02

OpenStudy (anonymous):

I'm just going to guide you. Working with the formula you provided, I'm going to rearrange it to make m the subject of the formula: \[K.E=\frac{ 1 }{ 2 } mv ^{2}\] Multiplying both sides by 2, we get: \[2 K.E = 2 \times(\frac{ 1 }{ 2 } mv ^{2} ) \] This then gives: \[2K.E= mv ^{2}\] Dividing both sides by v2, we get \[\frac{ 2 K.E }{ v ^{2} }= \frac{ mv ^{2} }{ v ^{2} }\] Cancelling out the two v2, we are left with: \[m=\frac{ 2 K.E }{ v ^{2} }\] That's the formula for finding the mass. All you have to do is plug in the values. K.E = 105 J and v= 10 m/s.

OpenStudy (anonymous):

Follow the formula and see if you'll get 2.1 kg.

OpenStudy (firejay5):

@kym02 I didn't change the original formula, I just used m = 2 X KE / v^2

OpenStudy (anonymous):

yes, you transposed! did you get back 2.1 kg?